After the RH was switched off in LN3 at 08:26 UTC, no changes were observed in the 3 Hz comb.
Since we were still in LN3, we agreed with Maria and Michal to try disconnecting the RH cables to check for any effects on the 3 Hz comb. Andrea went to the WE to perform the operation. The sequence of actions is listed below:
-
08:57 UTC – RH cables unplugged from the power supply (still powered, but output set to 0 V).
-
09:02 UTC – RH cables replugged.
As observed in other tests, the 50 Hz increased significantly (see entry 63151). However, some peaks of the 3 Hz comb disappeared or showed a reduced amplitude.
To confirm the results, we repeated the test:
-
09:11 UTC – RH cables unplugged again.
-
09:22 UTC – ITF unlocked.
The same behavior as in the previous data segment was observed:
-
Figure 1: Spectrogram covering the entire test duration.
-
Figure 2: Comparison between RH off but connected (purple curve) and RH off and disconnected (blue curve).
-
Figure 3: Zoomed-in view of Figure 2 for better detail.
-
Figure 4: spectrum highlighting the 50 Hz sidebands.
-
Figure 5: RH cable disconnected
After the unlock, Andrea reconnected the RH and left it in place.
Side note: the WE RH was already disconnected during the investigation of the 25 min glitches (63246).
To be investigated: RH power supply and its cabling